# Phase transitions for unique codings of fat Sierpinski gasket with multiple digits

### v1

@@ 5. Generalized Komornik-Loreti constant $\beta_{c}(M)$ for $M=3N+2$
[…]
@@
$x$$y$$\alpha_{00}$$\alpha_{10}$$\alpha_{20}$$\alpha_{01}$$\alpha_{02}$$\alpha_{11}$$\frac{1}{\beta}$$\frac{2}{\beta}$$\frac{2}{\beta-1}$$\frac{1}{\beta}$$\frac{2}{\beta}$$\frac{2}{\beta-1}$
@@
We also define the conjugacy of $(\gamma_{i})$ by
@@
To prove Theorem 1.4 for $M=3N+2$ we will construct a sequence of blocks $\left\{{\mathbf{b}}_{n}\right\}_{n=0}^{\infty}$ in ${\Omega}_{M}^{*}$. Let
$\pazocal A_{2}$ :=$:=\left\{\alpha_{(N+1)(N+1)}\alpha_{(N+1)N}\alpha_{N(N+1)},\quad\alpha_{N(N+1)}\alpha_{(N+1)N}\alpha_{(N+1)(N+1)},\right.$
@@
Then the sequence $\left\{{\mathbf{b}}_{n}\right\}_{n=0}^{\infty}\subset\pazocal A_{1}^{*}$ is defined as follows (see Figure 6):
@@
${\mathbf{b}}_{0}\sim\left(\begin{array}[]{ccc}N+1&N+1&N\\ N+1&N&N+1\\ N&N+1&N+1\end{array}\right)$
@@
for any $n\geq 0$.
The following result can be proved in a similar way as that for Proposition 4.11.
@@
So, in this case we don’t need a lemma like Lemma 4.5. All other properties are similar to that for $M=3N+1$. For example, we have a similar characterization on patterns appearing in sequences of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$.
@@
The proof of Proposition 5.3 is very similar to Proposition 4.7, and we leave it as an exercise for interested readers.
@@ 6. Generalized Komornik-Loreti constant for $\widetilde{{\mathbf{U}}}_{\beta,M}$ with $M=3N+3$
Let $M=3N+3$ with $N\in\mathbb{N}_{0}$. Recall that $\beta_{c}(M)$ is defined in Definition 1.2. In this section we will prove Theorem 1.4 for $M=3N+3$ that $\widetilde{{\mathbf{U}}}_{\beta,M}$ has positive Hausdorff dimension if and only if $\beta>\beta_{c}(M)$. Completely different from the case for $M=1$, we need more effort to find the critical base $\beta_{c}(M)$ for $M=3N+3$.
@@
[…]
@@
where $\sigma$ is the left shift map.
@@
Accordingly, we define the conjugacy of $(\eta_{i})$ by
@@
So, in the following we assume $10\leq k<2^{n}-$ 1. Note by (6.1) that
@@
where the second inequality follows by using the property of $(\tau_{i})$ that $\tau_{k+1}\ldots\tau_{2^{n}}\preccurlyeq\tau_{1}\ldots\tau_{2^{n}-k}$. This proves the second inequality in (6.4).
@@
So in the following we assume $10\leq k<2^{n}-$ 1. Then by (6.1) we have
@@
Case (III). $i=3k+2$ with $0\leq k<2^{n}$. Then $\eta_{i+1}=N+1+\tau_{k+1}\geq N+1>\hat{\eta}_{1}$, which gives the first inequality of (6.4). On the other hand, if $k=2^{n}-1$ then we have $\eta_{i+1}=N+1+\tau_{2^{n}}=N+2=\eta_{1}$; and if $10\leq k<2^{n}-$ 1, then by (6.1) we have
@@
Note that the Thue-Morse sequence $(\tau_{i})_{i=0}^{\infty}$ satisfies $\tau_{0}=0,\tau_{2k}=\tau_{k}$ and $\tau_{2k+1}=1-\tau_{k}$ for all $k\geq 0$. Then by Definition 1.2 and (6.1) it follows that $(\eta_{i})$ is an expansion of $1$ in base $\beta_{c}=\beta_{c}(M)$. So, by Lemmas 2.1 and 6.2 it follows that $\delta(\beta_{c})=\eta_{1}\eta_{2}\ldots$. Hence, by (6.1) we obtain
@@
Suppose on the contrary that $\beta_{c}\in(N+2,N+3)$ is algebraic. Then $\beta_{c}^{-3}\in(0,1)$ is also algebraic, and by Lemma 4.4 it follows that the left-hand side of the above equation is transcendental, while the right hand-side is algebraic, leading to a contradiction. So, $\beta_{c}$ is transcendental. ∎
@@ 6.1. Patterns in $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$
$x$$y$$\alpha_{00}$$\alpha_{10}$$\alpha_{20}$$\alpha_{30}$$\alpha_{01}$$\alpha_{02}$$\alpha_{03}$$\alpha_{11}$$\alpha_{21}$$\alpha_{12}$$\frac{1}{\beta}$$\frac{2}{\beta}$$\frac{3}{\beta}$$\frac{3}{\beta-1}$$\frac{1}{\beta}$$\frac{2}{\beta}$$\frac{3}{\beta}$$\frac{3}{\beta-1}$
@@
In the following we will study possible block patterns appearing in sequences of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. To describe these patterns we construct a sequence of blocks $\left\{{\mathbf{c}}_{n}\right\}_{n=0}^{\infty}$ in ${\Omega}_{M}^{*}$. Let
$\pazocal A_{3}$ :=$:=\left\{\alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+2)N},\quad\alpha_{N(N+1)}\alpha_{(N+1)(N+2)}\alpha_{(N+1)N},\right.$
@@
Then the sequence $\left\{{\mathbf{c}}_{n}\right\}_{n=0}^{\infty}\subset\pazocal A_{3}^{*}$ is define recursively by (see Figure 7)
@@
Next we define recursively a sequence of blocks $({\mathbf{v}}_{n})$ which converges to $(\eta_{i})$ componentwisely:
@@
Comparing the definitions of $({\mathbf{v}}_{n})$ and $({\mathbf{c}}_{n})$ in (6.11) and (6.10) respectively, it follows that
@@
for any $n\geq 0$.
@@
However, the forbidden block $(N+1)(N+2)(N+1)$ occurs in $\overline{d_{m+1}^{\oplus}}\ldots\overline{d_{m+9}^{\oplus}}$ in the first cases, and occurs in $d_{m+1}^{2}\ldots d_{m+9}^{2}$ in the second case, leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. So, $d_{m+4}=\alpha_{N(N+1)}$.
@@
Since the proof of (ii) is similar, here we only prove (i). Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. By the same argument as in the proof of Lemma 6.5 we may assume $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. Then by Lemma 6.4 there exists $m_{*}\in\mathbb{N}$ such that each component of $d_{m_{*}+j}$ belongs to $\left\{N,N+1,N+2\right\}$ for all $j\geq 1$. Take $m>m_{*}$. Suppose $d_{m}^{1}<M$ and $d_{m+1}\ldots d_{m+6}={\mathbf{c}}_{2}^{+}=\alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+2)(N+1)}\alpha_{N(N+1)}\alpha_{(N+1)(N+2)}\alpha_{(N+2)N}$ for some $m>m_{*}$. Then by (2.1), Propositions 2.3 and 6.1 it follows that $d_{m+7}\in\left\{\alpha_{N(N+1)},\alpha_{N(N+2)}\right\}$.
If $d_{m+7}=\alpha_{N(N+2)}$, then by (2.1), Propositions 2.3 and 6.1 we obtain
@@
Thus, the forbidden block $(N+1)(N+2)(N+1)$ appears in $\overline{d_{m+1}^{\oplus}}\ldots\overline{d_{m+9}^{\oplus}}$, which leads to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. So, $d_{m+7}=\alpha_{N(N+1)}$.
@@
which will be done by induction on $n$. If $n=0$ or $1$, then (6.14) holds by Lemmas 6.5 and 6.6. Now suppose (6.14) holds for all $n\leq k$ for some $k\in\mathbb{N}$, and we consider it for $n=k+1$.
@@
Note that $d_{m+1}\ldots d_{m+3\cdot 2^{k+1}}={\mathbf{c}}_{k+2}^{+}={\mathbf{c}}_{k+1}^{+}\Phi_{3}({\mathbf{c}}_{k}^{+}){\mathbf{c}}_{k}^{+}$. Then $d_{m+3\cdot(2^{k}+2^{k-1})}^{1}=N+1<M$ and $d_{m+3(2^{k}+2^{k-1})+1}\ldots d_{m+3\cdot 2^{k+1}}={\mathbf{c}}_{k}^{+}$. Then by the induction hypothesis for $n=k-1$ we have
@@
If it takes $\Phi_{3}({\mathbf{c}}_{k})$, then $d^{1}_{m+3\cdot 2^{k+1}+1}\ldots d^{1}_{m+3\cdot(2^{k+1}+2^{k-1})}=\phi_{3}({\mathbf{v}}_{k})\succ\phi_{3}({\mathbf{v}}_{k}^{+})$, leading to a contradiction with (6.15). So, $d_{m+3\cdot 2^{k+1}+1}\ldots d_{m+3\cdot(2^{k+1}+2^{k-1})}=\Phi_{3}({\mathbf{c}}_{k}^{+})$. Since $d_{m+3\cdot 2^{k+1}}^{1}=N+2>0$, by the induction hypothesis we obtain that $d_{m+3\cdot 2^{k+1}+1}\ldots d_{m+3\cdot(2^{k+1}+2^{k})}=\Phi_{3}({\mathbf{c}}_{k+1}^{+})$ or $\Phi_{3}({\mathbf{c}}_{k+1})$. In terms of (6.15) the block $\Phi_{3}({\mathbf{c}}_{k+1})$ is not allowed. So,

### v2

@@ 5. Generalized Komornik-Loreti constant $\beta_{c}(M)$ for $M=3N+2$
where $(\tau_{i})_{i=0}^{\infty}=0110100110010110\ldots\in\left\{0,1\right\}^{\mathbb{N}}$ is the classical Thue-Morse sequence. Then $(\gamma_{i})$ begins with (N+1)(N+1)(N+1)N(N+1)(N+1)N(N+1)N(N+1)(N+1)(N+1)(N+1)(N+1)(N+1)N(N+1)(N+1)N(N+1)N(N+1)$(N+1)$. Similar to Lemma 4.1, the sequence $(\gamma_{i})$ can also be obtained by substitutions. Let $\varphi_{2}:0\mapsto N(N+1)(N+1),\penalty 10000\ 1\mapsto(N+1)N(N+1).$ Then
@@
[…]
@@
We also define the conjugate of $(\gamma_{i})$ by
@@
To prove Theorem 1.4 for $M=3N+2$ we will construct a sequence of blocks $\left\{{\mathbf{b}}_{n}\right\}_{n=0}^{\infty}$ in ${\Omega}_{M}^{*}$. Similar to the words $(\mathbf{a}_{n})$ in the previous section, these words $({\mathbf{b}}_{n})$ are like basic bricks to describe the set $\widetilde{{\mathbf{U}}}_{\beta,M}$ for $\beta\leq\beta_{c}(M)$ with $M=3N+2$. Let
$\pazocal A_{2}$ $:=\left\{\alpha_{(N+1)(N+1)}\alpha_{(N+1)N}\alpha_{N(N+1)},\quad\alpha_{N(N+1)}\alpha_{(N+1)N}\alpha_{(N+1)(N+1)},\right$.
@@
Then the sequence $\left\{{\mathbf{b}}_{n}\right\}_{n=0}^{\infty}\subset\pazocal A_{2}^{*}$ is defined as follows (see Figure 6):
@@
${\mathbf{b}}_{0}\sim\left(\begin{array}[]{ccc}N+1&N+1&N\\ N+1&N&N+1\\ N&N+1&N+1\end{array}\right),$
@@
for any $n\geq 0$. Recall that for a block ${\mathbf{d}}=d_{1}d_{2}\ldots d_{k}\in\Omega_{M}^{*}$ we write ${\mathbf{d}}^{1}=d_{1}^{1}d_{2}^{1}\ldots d_{k}^{1},{\mathbf{d}}^{2}=d_{1}^{2}d_{2}^{2}\ldots d_{k}^{2}$ and $\overline{{\mathbf{d}}^{\oplus}}=\overline{d_{1}^{\oplus}}\,\overline{d_{2}^{\oplus}}\ldots\overline{d_{k}^{\oplus}}$, where each $d_{i}=(d_{i}^{1},d_{i}^{2})$ and $\overline{d_{i}^{\oplus}}=M-d_{i}^{\oplus}$ with $d_{i}^{\oplus}=d_{i}^{1}+d_{i}^{2}$.
The following result can be proved in a similar way to that for Proposition 4.12.
@@
So, in this case we don’t need a lemma like Lemma 4.5. All the other properties are similar to those for $M=3N+1$. For example, we have a similar characterization on patterns appearing in sequences of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$.
@@
The proof of Proposition 5.3 is very similar to that of Proposition 4.8, and we leave it as an exercise for interested readers.
@@ 6. Generalized Komornik-Loreti constant for $\widetilde{{\mathbf{U}}}_{\beta,M}$ with $M=3N+3$
Let $M=3N+3$ with $N\in\mathbb{N}_{0}$. Recall that $\beta_{c}(M)$ is defined in Definition 1.2. In this section we will prove Theorem 1.4 for $M=3N+3$, i.e., $\widetilde{{\mathbf{U}}}_{\beta,M}$ has positive Hausdorff dimension if and only if $\beta>\beta_{c}(M)$. Completely different from the case for $M=1$, we need more effort to find the critical base $\beta_{c}(M)$ for $M=3N+3$.
@@
Similar to Lemma 4.1, the sequence $(\eta_{i})$ can also be obtained by substitutions. Let $\varphi_{3}:0\mapsto(N+1)(N+2)N,\penalty 10000\ 1\mapsto(N+2)N(N+1).$ Then
@@
where $\sigma$ is the left-shift map.
@@
Accordingly, we define the conjugate of $(\eta_{i})$ by
@@
So, in the following we assume $0\leq k<2^{n}-1$. Note by (6.1) that
@@
where the second inequality follows by using the property of $(\tau_{i})$, i.e., $\tau_{k+1}\ldots\tau_{2^{n}}\preccurlyeq\tau_{1}\ldots\tau_{2^{n}-k}$. This proves the second inequality in (6.4).
@@
So in the following we assume $0\leq k<2^{n}-1$. Then by (6.1) we have
@@
Case (III). $i=3k+2$ with $0\leq k<2^{n}$. Then $\eta_{i+1}=N+1+\tau_{k+1}\geq N+1>\hat{\eta}_{1}$, which gives the first inequality of (6.4). On the other hand, if $k=2^{n}-1$ then we have $\eta_{i+1}=N+1+\tau_{2^{n}}=N+2=\eta_{1}$; and if $0\leq k<2^{n}-1$, then by (6.1) we have
@@
Note that the Thue-Morse sequence $(\tau_{i})_{i=0}^{\infty}$ satisfies $\tau_{0}=0,\tau_{2k}=\tau_{k}$ and $\tau_{2k+1}=1-\tau_{k}$ for all $k\geq 0$. Then by Definition 1.2 and (6.1) it follows that $(\eta_{i})$ is an expansion of $1$ in the base $\beta_{c}=\beta_{c}(M)$. So, by Lemmas 2.1 and 6.2 it follows that $\delta(\beta_{c})=\eta_{1}\eta_{2}\ldots$. Hence, by (6.1) we obtain
@@
Suppose on the contrary that $\beta_{c}\in(N+2,N+3)$ is algebraic. Then $\beta_{c}^{-3}\in(0,1)$ is also algebraic, and by Lemma 4.4 it follows that the left-hand side of the above equation is transcendental, while the right-hand side is algebraic, leading to a contradiction. So, $\beta_{c}$ is transcendental. ∎
@@ 6.1. Patterns in $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$
[…]
@@
In the following we will study possible block patterns appearing in sequences of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. To describe these patterns we construct a sequence of blocks $\left\{{\mathbf{c}}_{n}\right\}_{n=0}^{\infty}$ in ${\Omega}_{M}^{*}$. Similar to the words $(\mathbf{a}_{n})$ and $({\mathbf{b}}_{n})$ for the cases $M=3N+1$ and $M=3N+2$ respectively, these words $({\mathbf{c}}_{n})$ play an important role to describe $\widetilde{{\mathbf{U}}}_{\beta,M}$ for $\beta\leq\beta_{c}(M)$ with $M=3N+3$. Let
$\pazocal A_{3}$ $:=\left\{\alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+2)N},\quad\alpha_{N(N+1)}\alpha_{(N+1)(N+2)}\alpha_{(N+1)N},\right$.
@@
Then the sequence $\left\{{\mathbf{c}}_{n}\right\}_{n=0}^{\infty}\subset\pazocal A_{3}^{*}$ is defined recursively by (see Figure 7)
@@
Next we recursively define a sequence of blocks $({\mathbf{v}}_{n})$ which converges to $(\eta_{i})$ componentwisely:
@@
By comparing the definitions of $({\mathbf{v}}_{n})$ and $({\mathbf{c}}_{n})$ in (6.11) and (6.10) respectively, it follows that
@@
for any $n\geq 0$. Remember that for a block ${\mathbf{d}}=d_{1}d_{2}\ldots d_{k}\in\Omega_{M}^{*}$ we write ${\mathbf{d}}^{1}=d_{1}^{1}d_{2}^{1}\ldots d_{k}^{1},{\mathbf{d}}^{2}=d_{1}^{2}d_{2}^{2}\ldots d_{k}^{2}$ and $\overline{{\mathbf{d}}^{\oplus}}=\overline{d_{1}^{\oplus}}\,\overline{d_{2}^{\oplus}}\ldots\overline{d_{k}^{\oplus}}$, where each $d_{i}=(d_{i}^{1},d_{i}^{2})$ and $\overline{d_{i}^{\oplus}}=M-d_{i}^{\oplus}$ with $d_{i}^{\oplus}=d_{i}^{1}+d_{i}^{2}$.
@@
However, the forbidden block $(N+1)(N+2)(N+1)$ occurs in $\overline{d_{m+1}^{\oplus}}\ldots\overline{d_{m+9}^{\oplus}}$ in the first case, and occurs in $d_{m+1}^{2}\ldots d_{m+9}^{2}$ in the second case, leading to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. So, $d_{m+4}=\alpha_{N(N+1)}$.
@@
Since the proof of (ii) is similar, here we only prove (i). Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. By the same argument as in the proof of Lemma 6.5 we may assume $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. Then by Lemma 6.4 there exists $m_{*}\in\mathbb{N}$ such that each component of $d_{m_{*}+j}$ belongs to $\left\{N,N+1,N+2\right\}$ for all $j\geq 1$. Take $m>m_{*}$. Suppose $d_{m}^{1}<M$ and $d_{m+1}\ldots d_{m+6}={\mathbf{c}}_{2}^{+}=\alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+2)(N+1)}\alpha_{N(N+1)}\alpha_{(N+1)(N+2)}\alpha_{(N+2)N}$ for some $m>m_{*}$. Then by (2.1), and Propositions 2.3 and 6.1 it follows that $d_{m+7}\in\left\{\alpha_{N(N+1)},\alpha_{N(N+2)}\right\}$.
If $d_{m+7}=\alpha_{N(N+2)}$, then by (2.1), and Propositions 2.3 and 6.1 we obtain
@@
Thus, the forbidden block $(N+1)(N+2)(N+1)$ appears in $\overline{d_{m+1}^{\oplus}}\ldots\overline{d_{m+9}^{\oplus}}$, which leads to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. So, $d_{m+7}=\alpha_{N(N+1)}$.
@@
which will be done by induction on $n$. If $n=0$ or $1$, then (6.14) holds by Lemmas 6.5 and 6.6. Now suppose (6.14) holds for all $n\leq k$ with $k\in\mathbb{N}$, and we consider it for $n=k+1$.
@@
Note that $d_{m+1}\ldots d_{m+3\cdot 2^{k+1}}={\mathbf{c}}_{k+2}^{+}={\mathbf{c}}_{k+1}^{+}\Phi_{3}({\mathbf{c}}_{k}^{+}){\mathbf{c}}_{k}^{+}$. Then $d_{m+3\cdot(2^{k}+2^{k-1})}^{1}=N+1<M$ and $d_{m+3(2^{k}+2^{k-1})+1}\ldots d_{m+3\cdot 2^{k+1}}={\mathbf{c}}_{k}^{+}$. Then by the induction hypothesis for n=k1n=k-1 we have
@@
If it takes $\Phi_{3}({\mathbf{c}}_{k})$, then $d^{1}_{m+3\cdot 2^{k+1}+1}\ldots d^{1}_{m+3\cdot(2^{k+1}+2^{k-1})}=\phi_{3}({\mathbf{v}}_{k})\succ\phi_{3}({\mathbf{v}}_{k}^{+})$, leading to a contradiction to (6.15). So, $d_{m+3\cdot 2^{k+1}+1}\ldots d_{m+3\cdot(2^{k+1}+2^{k-1})}=\Phi_{3}({\mathbf{c}}_{k}^{+})$. Since $d_{m+3\cdot 2^{k+1}}^{1}=N+2>0$, by the induction hypothesis we obtain that $d_{m+3\cdot 2^{k+1}+1}\ldots d_{m+3\cdot(2^{k+1}+2^{k})}=\Phi_{3}({\mathbf{c}}_{k+1}^{+})$ or $\Phi_{3}({\mathbf{c}}_{k+1})$. In terms of (6.15) the block $\Phi_{3}({\mathbf{c}}_{k+1})$ is not allowed. So,